# Surrogate models for diffusion on graphs via sparse polynomials

### v2

@@ 4.2 Validation on real-world examples
To validate our theoretical results on real-world examples, we approximate $u_{2}$ (the solution of the diffusion equation at node 2) over datasets retrieved from Twitter and from Facebook, respectively. We find communities in each network using the fluid community detection algorithm [32].
@@ 4.2.1 The Twitter dataset
The Twitter network [33, 34] represents the Twitter interaction network for the 117th United States Congress, both House of Representatives and Senate. The base data was collected via the Twitter’s API, then the empirical transmission probabilities (represented by weighted edges) were quantified according to the fraction of times one member retweeted, quote tweeted, replied to, or mentioned another member’s tweet. The number of nodes of the graph is $|V|=475$ and the number of edges is $|E|=13,289$. We detected 2 communities exploiting the fluid community detection algorithm, which contain, respectively, 110 and 365 nodes.
@@
We computed the average RMSE on the approximation of $u_{2}$ over 5 runs, ranging the number of training sample points from 10 to 60 random sample points. The RMSE is computed over 1000 test random sample points.
The experiments substantially confirm the trend already observed in Section 4.1.1 of greater performance of least squares optimization, compared to the ones of QCBP, when the problem is overdetermined, and the ability of compressed sensing to recover good approximation in the underdetermined regime.
@@ 4.2.2 The Facebook dataset
This dataset consists of “circles” (or “friends lists”) from Facebook. Facebook data was collected from survey participants using a specific Facebook app [34, 35]. The size of this network is one order of magnitude larger than Twitter’s. The number of nodes of the graph is $|V|=4,039$ and the number of edges is $|E|=88,234$. We detected 2 communities employing the fluid community detection algorithm, which contain, respectively, 887 and 3152 nodes.
@@ 5 Conclusion
Looking ahead, we identify several directions for future work. First, we aim to extend our numerical results for diffusion on graphs to cases involving time-dependent diffusion, covering both synthetic and real-life examples. Additionally, our analysis could be broadened to explore alternative parameterizations of the problem, such as focusing on initial conditions rather than just the diffusion coefficient matrix.
Beyond the diffusion equation, it would be valuable to develop polynomial-based surrogate models for more complex dynamic models on graphs. A notable example is the Kuramoto model [36], which describes how groups of oscillators tend to synchronize based on their network structure.
The insights provided by sparse polynomial surrogate models may also contribute to the characterization of deep learning surrogate models through the emerging field of practical existence theorems [37, 38]. The theory developed in this work could help bridge the gap in the literature on practical existence theorems relating to the approximation of functions on graphs.
Finally, the features highlighted in our work regarding time-dependent graph diffusion mechanisms within communities offer valuable insights for modeling deep learning architectures, particularly Graph Neural Networks (GNNs) [39, 40]. In recent years, GNNs have evolved based on diffusion rules [41, 19, 42] that generalize the standard message-passing scheme [43]. This evolution aims to address known challenges in the graph learning community, including oversmoothing, oversquashing, and the information bottleneck problem. Sparse polynomial approximation can be leveraged in the context of graph diffusion to design innovative GNN architectures capable of capturing more complex dynamics in community-aware scenarios.
@@ Appendix A Supplementary information about multi-index sets
Definition 4 (Lower sets).
@@
Intuitively, lower sets do not have any “holes”. Note that the sets illustrated in Figure 1 are lower sets, and for example, $S=\{(0,0),(2,0)\}$ is not a lower set. Lower sets are discussed in more detail in, e.g., [1], but an important property is that a lower set $S$ of fixed size $s$ cannot contain multi-indices that are too far from the origin. In other words, this additional structure ensures that we obtain a finite set by taking the union of lower sets of cardinality less than or equal to $s$.
Remark 4.
The union of lower sets with cardinality $s$ is a hyperbolic cross $\Lambda_{s-1}^{\mathrm{HC}}$ of order s1s-1 (see [1, Proposition 2.5]). This emphasizes the connection between the hyperbolic cross index set and lower sets, and the importance of the hyperbolic cross in relation to sparse polynomial approximation, when the set $S$ is unknown.
@@ Appendix B Proof of Corollary 2
To prove Corollary 2 we need an auxiliary result. Specifically, we need an adaptation of [26, Theorem 6.1] to the finite-dimensional case.
Theorem 2.
Let $0<\epsilon<1$, $\varrho$ be either the uniform or Chebyshev measure on $\mathcal{U}=[-1,1]^{d}$, $m\geq 3$ and $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{m}\sim_{\text{i.i.d}}\varrho$. Then, there exists a set $S\subseteq\mathbb{N}_{0}^{d}$ of cardinality $|S|\leq\lceil m/\log(m/\epsilon)\rceil$ such that the following holds with probability at least $1-\epsilon$ for each fixed $f:\mathcal{U}\rightarrow\mathbb{C}$ that admits a holomorphic extension to a Bernstein polyellipse $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathbb{C}^{d}$. For any $\boldsymbol{n}\in\mathbb{C}^{m}$, the approximation $\hat{f}$ to the problem (9) is unique and satisfies, for every $0<p<1$, the bound $\|f-\hat{f}\|_{L_{\varrho}^{2}(\mathcal{U})}\leq C(\boldsymbol{\rho},p)\cdot(m/\log(m/\epsilon))^{\frac{1}{2}-\frac{1}{p}}+2\cdot\|\boldsymbol{n}\|_{\infty}$.
Proof.
To prove the theorem it is enough to apply [26, Theorem 6.1] combined with [26, Remark 5.2]. We only need to show that the support $S$ of the best $s$-term approximation to $f$ satisfies $S\subset\mathbb{N}_{0}^{d}\times\{0\}^{\mathbb{N}}$. Let $\mathcal{F}=\{\boldsymbol{\nu}\in\mathbb{N}_{0}^{\mathbb{N}}:\|\boldsymbol{\nu}\|_{0}<\infty\}\subset\mathbb{N}_{0}^{\mathbb{N}}$.
Recall that $f(\boldsymbol{x})=f(x_{1},\ldots,x_{d})$. Let $\boldsymbol{\nu}\in\mathcal{F}$ such that $\nu_{k}\neq 0$ for some $k>d$. Then
@@
Hence, the best $s$-term approximation support $S$ can be chosen in $\mathbb{N}_{0}^{d}\times\{0\}^{\mathbb{N}}$. ∎
The proof of Corollary 2 comes directly by combining Theorem 2 with Theorem 1 considering the following setting: $\mathcal{H}=\mathcal{E}_{\boldsymbol{\rho}}$, $\mathcal{K}=\mathcal{E}_{\boldsymbol{\rho}^{\prime}}$, where $\boldsymbol{\rho}^{\prime}>\boldsymbol{\rho}$ (strict inequality applies to every component) and, consequently, $\mathcal{O}=\mathring{\mathcal{K}}$.

### v3

@@ 4.1.4 Impact of number of nodes per community for fixed dimension
4.2 Synthetic graphs (time-and-edge-dependent case)
The aim of this section is to assess the performance of sparse polynomial surrogates over a time-dependent version of the numerical study of Section 4.1.1, reproducing the experiments in Figures 2 and 3 in the time-and-edge-dependent case. In particular, we consider a diffusivity coefficient matrix $\mathbf{C}(t,\boldsymbol{y})$ as in equation (18) where
$h_{\sigma(1,1)}(t)=h_{1}(t):=\frac{1}{2}\left(1+\frac{t}{T}\right),$
$h_{\sigma(1,2)}(t)=h_{\sigma(2,1)}(t)=h_{2}(t):=\frac{1}{3}(2+\cos(t)),$
$h_{\sigma(2,2)}(t)=h_{3}(t):=\frac{1}{3}(2+\sin(t)).$
These functions have been chosen in such a way to take strictly positive value less than or equal to $1$ for $t\in[0,T]$.
We start by studying the polynomial coefficients and the best $s$-term approximation of the function $f_{v}(\boldsymbol{y})$. Figure 11 shows the results of the same experiment as in Figure 2 for the time-and-edge-dependent problem considered. Analogously to the edge-dependent case, we observe a clear super-algebraic decay of the sorted coefficients and of the best $s$-term approximation error. This numerically validates the holomorphic regularity of the underlying function established by Theorem 3.1. In addition, we conduct the same experiment as in Figure 3 in the time-and-edge-dependent case to study the convergence of the proposed surrogate models by plotting the average RMSE as a function of the number of samples $m$. The results, reported in Figure 12, show a convergence behavior very similar to that observed in the edge-dependent case.
Figure 11: Time-and-edge-dependent case. Absolute polynomial coefficients of the parametric map $f_{v}(\boldsymbol{y})$ (left: unsorted, center: sorted) and best $s$-term approximation error (right).
Figure 12: Time-and-edge-dependent case. Convergence plot of average RMSE vs. number of sample points for $d=3$.
In summary, a comparison of the results in Figures 11 and 12 with those in Figures 2 and 3 shows that the introduction of a time-dependent diffusivity matrix does not numerically affect the performance of the surrogate method in the example considered.
@@ 4.3 Validation on real-world examples
To validate our theoretical results on real-world examples, we approximate $u_{2}$ (the solution of the diffusion equation at node 2) over datasets retrieved from Twitter and from Facebook, respectively. We find communities in each network using the fluid community detection algorithm pares2018fluid .
@@ 4.3.1 The Twitter dataset
The Twitter network fink2023twitter ,snapnets represents the Twitter interaction network for the 117th United States Congress, both House of Representatives and Senate. The base data was collected via the Twitter’s API, then the empirical transmission probabilities (represented by weighted edges) were quantified according to the fraction of times one member retweeted, quote tweeted, replied to, or mentioned another member’s tweet. The number of nodes of the graph is $|V|=475$ and the number of edges is $|E|=13,289$. We detected 2 communities exploiting the fluid community detection algorithm, which contain, respectively, 110 and 365 nodes.
@@
We computed the average RMSE on the approximation of $u_{2}$ over 5 runs, ranging the number of training sample points from 10 to 60 random sample points. The RMSE is computed over 1000 test random sample points. The experiments substantially confirm the trend already observed in Section 4.1.1 of greater performance of least squares optimization, compared to the ones of QCBP, when the problem is overdetermined, and the ability of compressed sensing to recover good approximation in the underdetermined regime.
@@ 4.3.2 The Facebook dataset
This dataset consists of “circles” (or “friends lists”) from Facebook. Facebook data was collected from survey participants using a specific Facebook app snapnets leskovec2012learning . The size of this network is one order of magnitude larger than Twitter’s. The number of nodes of the graph is $|V|=4,039$ and the number of edges is $|E|=88,234$. We detected 2 communities employing the fluid community detection algorithm, which contain, respectively, 887 and 3152 nodes.
@@
4.4 Computational cost analysis
In this section we conduct an experimental analysis of the computational cost of the proposed surrogate modeling strategy in order to showcase its applicability. Here we consider the time-and-edge-dependent problem of Section 4.2 and the real-world graphs of Section 4.3. We denote as $t_{\mathrm{full}}$ the full order model time, i.e., the time needed to collect one sample by solving a diffusion equation. The polynomial surrogate has an offline and an online cost. The offline cost can be split into two contributions: namely, the cost of collecting $m$ samples—given by $m\cdot t_{\mathrm{full}}$—and the time $t_{\mathrm{coeff}}(\Lambda,m)$ needed to compute the polynomial coefficients. Finally, we denote as $t_{\mathrm{online}}(\Lambda)$ the online time for the surrogate model, i.e., the time needed to evaluate the compute polynomial at one point. In Figure 14 (left) we plot the average times of the offline and online cost of both surrogate and full order models (i.e., $t_{\mathrm{full}}$, $t_{\mathrm{coeff}}(\Lambda,m)$, and $t_{\mathrm{online}}(\Lambda)$) as a function of the number of samples $m$. Geometric means are computed over different random seeds: namely, 20 different seeds for the time-and-edge-dependent experiments on SBM, and 5 seeds for the edge-dependent case on real-world datasets.
Figure 14: Left: Comparison of the average computational cost for different case studies. Solid lines refer to the cost $t_{\mathrm{coeff}}(\Lambda,m)$ of computing the coefficients of the surrogate model. The dashed lines represent the evaluation time $t_{\mathrm{online}}(\Lambda)$ of the surrogate at a single parameter vector. Solid lines with star markers refer to the evaluation time $t_{\mathrm{full}}$ of the full order model. Right: Plot of $k_{\min}-m$ (see (28)) against the number of samples $m$. Top: time-and-edge-dependent diffusion on the SBM model. Center: edge-dependent diffusion on the Twitter dataset. Bottom: edge-dependent diffusion on the Facebook dataset.
In all cases, the online evaluation of the sparse polynomial surrogate is always faster than the computation of the full order model simulation by several order of magnitudes (i.e., $t_{\mathrm{online}}(\Lambda)\ll t_{\mathrm{full}}$). Moreover, while for real-world datasets the full order model evaluation of one sample takes more time even than the offline computation of the coefficients of the surrogates (which makes the latter much better performing), for the time-dependent case the situation is different. In fact, we can see that computing the coefficients for compressed sensing surrogates with hyperbolic cross index set with order $s=20$ introduces a trade off, in terms of how many evaluations would be needed to make the computation of the surrogate worth to replace the full order model. For this reason, we compute the minimum number of evaluations $k$ for which constructing the surrogate and evaluating it $k$ times is faster than evaluating the full order model $k$ times. This is defined by the inequality
[…]
which is solved by
$k\geq\frac{m\cdot t_{\mathrm{full}}+t_{\mathrm{coeff}}(\Lambda,m)}{t_{\mathrm{full}}-t_{\mathrm{online}}(\Lambda)}=:k_{\min}.$ (28)
In Figure 14 (right) we plot the difference $k_{\min}-m$ as a function of $m$. We plot $k_{\min}-m$ and not simply $k_{\min}$ to optimize visualization since in all cases, and especially for real-world datasets, we see that $k_{\min}\approx m$. Here $k_{\min}$ is averaged over multiple seeds as explained above.
@@ 5 Conclusion
Looking ahead, we identify several directions for future work. First, we aim to extend our numerical results for diffusion on graphs to cases involving time-dependent diffusion on further real-world graphs beyond the Twitter and Facebook datasets considered in this paper. In particular, we anticipate the benefit of the proposed surrogate modeling strategy to be most effective for very large-scale graphs. Additionally, our analysis could be broadened to explore alternative parametrizations of the problem, such as focusing on initial conditions rather than just the diffusion coefficient matrix.
Beyond the diffusion equation, it would be valuable to develop polynomial-based surrogate models for more complex dynamic models on graphs. A notable example is the Kuramoto model rodrigues2016kuramoto , which describes how groups of oscillators tend to synchronize based on their network structure.
The insights provided by sparse polynomial surrogate models may also contribute to the characterization of deep learning surrogate models through the emerging field of practical existence theorems adcock2022deep ; adcock2025near . The theory developed in this work could help bridge the gap in the literature on practical existence theorems relating to the approximation of functions on graphs.
Finally, the features highlighted in our work regarding time-dependent graph diffusion mechanisms within communities offer valuable insights for modeling deep learning architectures, particularly Graph Neural Networks (GNNs) scarselli2008graph ; kipf2016semi . In recent years, GNNs have evolved based on diffusion rules chamberlain2021beltrami ; chamberlain2021grand ; toppingunderstanding that generalize the standard message-passing scheme gilmer2017neural . This evolution aims to address known challenges in the graph learning community, including oversmoothing, oversquashing, and the information bottleneck problem. Sparse polynomial approximation can be leveraged in the context of graph diffusion to design innovative GNN architectures capable of capturing more complex dynamics in community-aware scenarios.
@@ Appendix A Supplementary information about multi-index sets
Definition 4 (Lower sets)
@@
Intuitively, lower sets do not have any “holes”. Note that the sets illustrated in Figure 1 are lower sets, and for example, $S=\{(0,0),(2,0)\}$ is not a lower set. Lower sets are discussed in more detail in, e.g., sparsepoly , but an important property is that a lower set $S$ of fixed size $s$ cannot contain multi-indices that are too far from the origin. In other words, this additional structure ensures that we obtain a finite set by taking the union of lower sets of cardinality less than or equal to $s$.
Remark 6
The union of lower sets with cardinality $s$ is a hyperbolic cross $\Lambda_{s-1}^{\mathrm{HC}}$ of order s1s-1 (see (sparsepoly, , Proposition 2.5)). This emphasizes the connection between the hyperbolic cross index set and lower sets, and the importance of the hyperbolic cross in relation to sparse polynomial approximation, when the set $S$ is unknown.
@@ Appendix B Proof of Corollary 2
To prove Corollary 2 we need an auxiliary result. Specifically, we need an adaptation of (adcock2022monte, , Theorem 6.1) to the finite-dimensional case.
Theorem B.1
Let $0<\epsilon<1$, $\varrho$ be either the uniform or Chebyshev measure on $\mathcal{U}=[-1,1]^{d}$, $m\geq 3$ and $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{m}\sim_{\text{i.i.d}}\varrho$. Then, there exists a set $S\subseteq\mathbb{N}_{0}^{d}$ of cardinality $|S|\leq\lceil m/\log(m/\epsilon)\rceil$ such that the following holds with probability at least $1-\epsilon$ for each fixed $f:\mathcal{U}\rightarrow\mathbb{C}$ that admits a holomorphic extension $\tilde{f}$ to a Bernstein polyellipse $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathbb{C}^{d}$. For any $\boldsymbol{n}\in\mathbb{C}^{m}$, the approximation $\hat{f}$ to the problem (9) is unique and satisfies, for every $0<p<1$, the bound $\|f-\hat{f}\|_{L_{\varrho}^{2}(\mathcal{U})}\leq C(\boldsymbol{\rho},p)\cdot\|\tilde{f}\|_{L^{\infty}(\mathcal{E}_{\boldsymbol{\rho}})}\cdot(m/\log(m/\epsilon))^{\frac{1}{2}-\frac{1}{p}}+2\cdot\|\boldsymbol{n}\|_{\infty}$.
Proof
To prove the theorem it is enough to apply (adcock2022monte, , Theorem 6.1) combined with (adcock2022monte, , Remark 5.2). We only need to show that the support $S$ of the best $s$-term approximation to $f$ satisfies $S\subset\mathbb{N}_{0}^{d}\times\{0\}^{\mathbb{N}}$. Let $\mathcal{F}=\{\boldsymbol{\nu}\in\mathbb{N}_{0}^{\mathbb{N}}:\|\boldsymbol{\nu}\|_{0}<\infty\}\subset\mathbb{N}_{0}^{\mathbb{N}}$. Recall that $f(\boldsymbol{x})=f(x_{1},\ldots,x_{d})$. Let $\boldsymbol{\nu}\in\mathcal{F}$ such that $\nu_{k}\neq 0$ for some $k>d$. Then
@@
Hence, the best $s$-term approximation support $S$ can be chosen in $\mathbb{N}_{0}^{d}\times\{0\}^{\mathbb{N}}$.
The proof of Corollary 2 comes directly by combining Theorem B.1 with Theorem 3.1 considering the following setting: $\mathcal{H}=\mathcal{E}_{\boldsymbol{\rho}}$, $\mathcal{K}=\mathcal{E}_{\boldsymbol{\rho}^{\prime}}$, where $\boldsymbol{\rho}^{\prime}>\boldsymbol{\rho}$ (strict inequality applies to every component) and, consequently, $\mathcal{O}=\mathring{\mathcal{K}}$.